A uniform metallic rod of length 1m is rigidly clamped at two points C(10 cm from one end) and D(30cm from other other end) and longitudinal waves are set up in the rod . Find the frequency of first overtone of longitudinal vibration the rod. 
 
γ(gamma of rod)= 1.6x10^11 N/m^2 Density of rod = 2500kg/m^3

Grade:11

1 Answers

Abhishek Kumar
9 years ago
Frequency of first overtone = 2V/2L =V/L
L = 20 cm = 0.2 m
V=\sqrt{Y/\rho }